OpenStudy (anonymous):

The altitude of an equilateral triangle is 7 root 3 units long. The length of one side of the triangle is ____ units. A. 7 B. 14 C. 14 root 3 D. (7 root 3) /2

OpenStudy (ash2326):

@angelicapad All sides are equal for an equilateral triangle, altitude for an equilateral triangle is also the median. So altitude will the half the side. See the diagram all sides are equal let, them be a and we have the altitude as \(7\sqrt 3\) In triangle ABD AB=a AD=\(7\sqrt 3\) BD=a/2 ABD is a right triangle, can you find a now?